# Detect Anything 3D in the Wild

### v2

@@ 7.1 Camera and Depth Module Details
This dense representation of the camera rays undergoes Laplace Spherical Harmonic Encoding (SHE) [54] to produce the embeddings $\mathbf{C}$. These embeddings are then passed to the depth module using the cross-attention mechanism.
@@ 8 Target-aware Metrics
In our work, we evaluate both traditional metrics and the target-aware metrics proposed by OVMono3D [74]. Under the target-aware paradigm, rather than prompting the model with all possible classes from an entire dataset, we only prompt it with the classes present in the current image during inference. This is designed to address two key challenges encountered:
@@
As illustrated in Figure 9, these issues are especially pronounced in the 3RScan [65] dataset. The left side shows the official 3RScan annotations, while the right side shows detections from Grounding DINO, which are largely misaligned with the dataset’s labeling conventions. Consequently, traditional evaluation metrics may yield misleading or inconsistent results, whereas target-aware metrics help mitigate these mismatches by restricting the evaluated classes to those actually present in the scene.

### v3

@@ 7.1 Camera and Depth Module Details
This dense representation of the camera rays undergoes Laplace Spherical Harmonic Encoding (SHE) [piccinelli2024unidepth] to produce the embeddings $\mathbf{C}$. These embeddings are then passed to the depth module using the cross-attention mechanism.
@@ 8 Target-aware Metrics
In our work, we evaluate both traditional metrics and the target-aware metrics proposed by OVMono3D [yao2024open]. Under the target-aware paradigm, rather than prompting the model with all possible classes from an entire dataset, we only prompt it with the classes present in the current image during inference. This is designed to address two key challenges encountered:
@@
As illustrated in Figure 9, these issues are especially pronounced in the 3RScan [wald2019rio] dataset. The left side shows the official 3RScan annotations, while the right side shows detections from Grounding DINO, which are largely misaligned with the dataset’s labeling conventions. Consequently, traditional evaluation metrics may yield misleading or inconsistent results, whereas target-aware metrics help mitigate these mismatches by restricting the evaluated classes to those actually present in the scene.
@@ 9.2 Ablation on Different Backbones
9.3 Ablation on DA3D Dataset
We ablate the impact of the DA3D dataset in Tab. 6. The additional data in DA3D primarily improves generalization to novel cameras, as Omni3D contains only two distinctive intrinsics for outdoor scenes.
Table 6: Ablation on training datasets. Unless specified, all models are trained on the Omni3D dataset. For the in-domain setting, prompts are provided by Cube R-CNN, while prompts for novel classes and novel datasets are generated by Grounding DINO.
Method In-domain Novel Class Novel Camera
$\rm AP^{omni3d}_{3D}$} $\rm{AP^{kit}_{3D}}$ $\rm{AP^{sun}_{3D}}$ $\rm{AP^{city}_{3D}}$ $\rm{AP^{3rs}_{3D}}$
Cube R-CNN 23.26 - - 8.22 / - -
OVMono3D 22.98 4.71 / 4.71 4.07 / 16.78 5.88 / 10.98 0.37 / 8.48
DetAny3D 24.33 23.75 / 23.75 7.63 / 20.87 8.31 / 11.68 0.64 / 9.56
$\rm DetAny3D_{DA3D}$ 24.92 25.73 / 25.73 7.63 / 21.07 11.05 / 15.71 0.65 / 9.58
9.4 Ablation on Inference Speed
We compare the inference speed of DetAny3D with prior methods in Table 7. DetAny3D runs at 1.5 FPS on a single KITTI image, which is slower than Cube R-CNN (33.3 FPS) and OVMono3D (7.1 FPS). This is a trade-off for stronger generalization across novel categories and cameras, as DetAny3D is designed as a foundation model rather than for real-time deployment.
Table 7: Inference speed comparison on KITTI.
Method Cube R-CNN OVMono3D DetAny3D
FPS $\uparrow$ 33.3 7.1 1.5
9.5 Per-category Performance on Novel Classes
[…]
Category OVMono3D DetAny3D
Board 4.83 6.02
Printer 16.23 60.22
Painting 2.80 5.11
Microwave 30.31 57.21
Tray 10.11 6.70
Podium 48.37 73.65
Cart 47.31 33.46
Tram 4.71 27.90
Easy Categories 20.58 33.79
Monitor 9.44 15.95
Bag 15.61 17.69
Dresser 29.08 41.75
Keyboard 9.13 9.52
Drawers 43.04 40.80
Computer 7.44 12.37
Kitchen Pan 9.98 8.70
Potted Plant 6.66 26.34
Tissues 12.45 12.95
Rack 10.21 9.04
Toys 5.24 16.14
Phone 3.89 4.42
Soundsystem 13.22 6.21
Fireplace 13.16 30.75
Hard Categories 13.47 18.05
All Categories 16.05 23.77
10 Limitations
Text Prompt Process. Our method leverages open-vocabulary 2D detectors such as Grounding DINO to convert text prompts into 2D box prompts. While effective, this strategy may cause semantic loss, as textual nuances are not directly injected into the 3D detection pipeline. Moreover, 2D detectors are known to perform poorly under heavy occlusion or partial visibility, introducing a domain gap when transferring their outputs to 3D tasks.
Inference Efficiency. Although DetAny3D achieves strong generalization across novel categories and camera settings, its inference speed (1.5 FPS) is significantly slower than existing lightweight 3D detectors. This limits its applicability in latency-sensitive scenarios such as real-time robotics or autonomous driving.
Lack of Temporal Modeling. Our current design operates on single-frame inputs and does not utilize temporal information from video sequences. Incorporating motion cues and enforcing temporal consistency could potentially improve detection accuracy and enable better integration into downstream video-based tasks, such as video knowledge distillation and temporal grounding.